OpenStudy (anonymous):

One press can run the day’s newspapers in 6 hours, while another can do the same job in 8 hours. After running together for 2 hours to complete the job, the faster one breaks. How much longer must the slower press run to finish the newspapers? see if you can solve it(:

hero (hero):

Sat you can help her

OpenStudy (anonymous):

who can get it done the fastessttttt

hero (hero):

Sat is faster than me so I let him do it

OpenStudy (anonymous):

hahah mkay dont be a sore loser !

hero (hero):

Nobody is even talking about that.

hero (hero):

Don't be immature

OpenStudy (anonymous):

combined rate is \[\frac{1}{8}+\frac{1}{6}=\frac{7}{24}\] and it ran for 2 hours so it did \[\frac{7}{24}\times 2=\frac{14}{24}\] of the job, leaving \[\frac{10}{24}\] of the job left. now only the slower one is working, so we need to solve \[\frac{1}{8} T=\frac{10}{24}=\frac{5}{12}\] for T we get \[T=\frac{40}{12}=\frac{10}{3}\]

OpenStudy (anonymous):

sat is a genius...

OpenStudy (anonymous):

so what is the answer buddy?

OpenStudy (anonymous):

so 3 hours and twenty minutes. then again i could have made a mistake.

OpenStudy (anonymous):

YAYYYYY your the pellet!

OpenStudy (anonymous):

more math?!